# KiDS-SBI: Simulation-Based Inference Analysis of KiDS-1000 Cosmic Shear

### v1

@@ 2 KiDS-1000 data
The Kilo-Degree Survey (KiDS) is a large public galaxy survey conducted by the European Southern Observatory using the OmegaCAM CCD mosaic camera (Kuijken, 2011) which is attached to the 2.6 m VLT Survey Telescope (VST). The survey covers approximately 1,350 deg2 between two distinct fields known as KiDS North and South which are fields located in the Northern and Southern Galactic Caps (with DEC $\approx$ 0º and DEC $\approx$ -30º, respectively). This area is the same as the one covered by the VISTA Kilo-degree INfrared Galaxy survey (VIKING; Edge et al. 2013), which means that both surveys together observe every object with a total of nine photometric bands: $ugriZYJHK_{\rm s}$. The analysis presented in this paper makes use of DR4 (Kuijken et al., 2019), also known as KiDS-1000, which covers approximately 1,000 deg2. The effective area covered by KiDS-1000 shape measurements shown in Fig. 1 extends over 773.3 deg2 (Joachimi et al., 2021). This area is calculated directly from the mosaic mask defined a the native OmegaCAM pixel scale of 0.213 arcsec.
@@
KiDS images are processed using Astro-WISE (McFarland et al., 2013) and the methodology outlined in Wright et al. (2019) to join the VST and VIKING photometric measurements. To obtain the shape measurements, Theli (Erben et al., 2005) is used to process the $r$-band images, while the gravitational shear estimates are then obtained from lensfit (Miller et al., 2007, 2013; Fenech Conti et al., 2017). Depending on the shape noise variance and the ellipticity measurement noise variance, lensfit assigns a weight to the shear measurement of each galaxy, $w_{i}$, which scales the shear signal such that the signal-to-noise ratio is optimal. Although KiDS-SBI currently does not model lensfit weights (all galaxies are weighted equally by default), the weights are carried through the measurement pipeline when measuring the cosmic shear summary statistics from the real KiDS-1000 data. If the lensfit weight is mischaracterised, this may lead to shear bias. For this purpose, the shape measurements in the KiDS-1000 gold sample are calibrated to determine any shear biases due to selection biases, noise, weight bias, point-spread function residuals or otherwise are calibrated (Giblin et al., 2021). The resulting estimate of the linear multiplicative shear bias per tomographic bin is then used to unbias the galaxy shear measurements in KiDS-1000. All details on the reduction of the images and the shape calibration are given in Kuijken et al. (2019); Wright et al. (2019); Giblin et al. (2021).
@@
The photometric redshifts are estimated from the KiDS 9-band photometry using Bayesian template-fitting as incorporated within the BPZ code (Benítez, 2000; Wright et al., 2019, 2020). The photometric redshifts are calibrated with spectroscopic redshift measurements made of galaxies in the KiDS sample (Hildebrandt et al., 2021) based on a subsample of $99\%$ completeness. The mapping from photometric to spectroscopic redshift, $P(z|z_{\mathrm{ph}})$, is learnt through a self-organising map (Wright et al., 2020; Hildebrandt et al., 2021). As is the case for previous KiDS-1000 analyses, we use the KiDS-1000 ”gold-sample” which only considers galaxies with reliable redshift and shape measurements. The maximum posterior photometric redshift is subsequently used to split the source galaxy catalogue into five tomographic bins at the following boundaries: $(0.1,0.3,0.5,0.7,0.9,1.2]$. These then give the redshift distributions shown in Fig. 2. As the cosmological constraints are sensitive to shifts in the redshift distributions, the KiDS-1000 analysis also allows the shift in the mean of each tomographic bin to vary freely in the cosmological inference as a nuisance parameter while incorporating for the correlations between them (Asgari et al., 2021; Heymans et al., 2021; van den Busch et al., 2022). This not only avoids biases in the cosmological parameter estimates but also provides an additional test on the redshift calibration.
@@ 3 Forward simulations
The forward simulations for cosmic shear analyses described in this paper are fully implemented and available in the KiDS-SBI [^7] Kilo-Degree Survey – Simulation-Based Inference; https://github.com/mwiet/kids_sbi module. This code is built within the Cosmological Survey Inference System (CosmoSIS [^8] https://github.com/joezuntz/cosmosis; Zuntz et al. 2015) and it is based upon the KiDS Cosmology Analysis Pipeline (KCAP [^9] https://github.com/KiDS-WL/kcap; Joachimi et al. 2021; Asgari et al. 2021; Heymans et al. 2021; Tröster et al. 2021). In addition, the KiDS-SBI pipeline is generalisable to other cosmic shear analyses.
Cosmological parameters, $\mathbf{\Theta}$ 3D matter power spectrum, $P_{\delta,\mathrm{nl}}(\mathbf{k},z)$: Sect. 3.1 Baryonic feedback Split space into concentric shells: $i,j...$: Sect. 3.2 Resolution along the line-of-sight Non-Limber 2D projection: Sect. 3.3 2D power spectra, $C_{\delta\delta}^{(ij)}(\ell)$ Generate log-normal random fields: Sect. 3.4 Matter, $\delta^{(i)}(\bm{\theta})$, & convergence fields, $\kappa^{(i)}(\bm{\theta})$ Intrinsic alignments, $\kappa_{\mathrm{IA}}^{(i)}(\bm{\theta})$: Sect. 3.5 Sample galaxies: Sect. 3.6 &\&& 3.7 Angular galaxy positions, $\bm{\theta}$, redshifts, $z_{\mathrm{true}}$, observed tomographic bin, $p$, and observed ellipticity, $\epsilon$ Survey characteristics Mask Redshift distributions Shear biases Point spread function Galaxy density Intrinsic galaxy shapes Depth variability: Sect. 3.8 Measure and correct shear: Sect. 4.1 Observed shear, $\epsilon^{\mathrm{corr}}$ Measure pseudo-Cls, subtract noise and binning: Sect. 4.2 Pseudo-Cl, $\tilde{C}_{\epsilon\epsilon}^{(pq)}(\ell)$
@@
The overall outline of the forward simulations is shown in Fig. 3. Firstly, the simulation is fed a set of cosmological and astrophysical parameters, $\mathbf{\Theta}$, from a sampler of our choice. All of these parameters are passed into CAMB (Lewis et al., 2000; Lewis & Challinor, 2002; Howlett et al., 2012) which calculates the three-dimensional matter power spectrum, $P_{\delta,\mathrm{nl}}$, at a given cosmology (see also Sect. 3.1 for more details). We incorporate baryonic feedback using HMCode (Mead et al., 2016). Next, the simulation splits the light-cone along the line-of-sight into concentric spherical shells centred at the observer, as required by our large-scale structure simulator of choice, GLASS (Tessore et al., 2023). Such an approach is computationally efficient as all observables are simulated with a coarse resolution along the line-of-sight. Within each shell, we sample log-normal random matter fields to efficiently model the two-point statistics with high accuracy, while also incorporating some higher-order fluctuations. As an added benefit, this simplifies the implementation of effects such as redshift evolution as well as survey characteristics, as they usually also follow a spherical geometry. On the other hand, the concentric shells also allow to split the Universe into shells, which are of a scale large enough such that a log-normal field becomes an accurate description of the matter over-densities within it (see Sects. 3.2 for details).
@@
Once we know the geometry of our matter shells, we can project the three-dimensional matter power spectrum using the redshift distributions of these shells in order to obtain their two-dimensional angular power spectra, $C_{\delta\delta}$. The latter describes the correlations between the matter over-densities in each shell. Since we also require accurate cross-correlations between different shells to compute the log-normal fields, the Limber approximation is not applicable here, as the cross-correlations are zero in that approximation with the shells being non-overlapping along the line-of-sight. To avoid this, we project the three-dimensional matter power spectrum with a non-Limber integral. This is usually an expensive computation, so we show here a novel implementation of the Levin method of integration (Levin, 1996) which allows for the efficient integration of the spherical Bessel functions found in non-Limber integrals (more details are given in Sect. 3.3).
@@
We can then construct log-normal matter fields which are consistent with these correlations and their associated cosmology. To do this, we use the Generator for Large Scale Structure (GLASS [^10] Generator for Large Scale Structure; https://github.com/glass-dev/glass, Tessore et al. 2023), which efficiently generates correlated log-normal random matter fields, $\delta$, for each shell. Then, it accurately computes the respective convergence field, $\kappa$, (see Sect. 3.4) while also incorporating an effective convergence field due to intrinsic galaxy alignments, $\kappa_{\mathrm{IA}}$ (see Sect. 3.5).
@@
Upon the construction of the matter, $\delta$, and convergence fields, $\kappa$, we use them to Poisson-sample galaxies that reside within the matter fields as well as their properties, i.e. intrinsic ellipticities and shear. During this step, with the aid of the SALMO module (Joachimi et al., 2021), we also consider survey characteristics, such as the survey footprint, its spatial variability in depth, in its redshift distributions, as well as in the shape and shot noise as described in Sect. 3.6. Thus, we obtain a full galaxy catalogue with shape information consistent with the input cosmology as well as with the previously mentioned survey properties.
@@
The last steps of the forward simulation pipeline involve the post-processing and compression of the simulated data catalogue. In principle, any data vector could be obtained here, such as two-point shear statistics (Asgari et al., 2021; van den Busch et al., 2022), the shear fields (Porqueres et al., 2022; Loureiro et al., 2023) or even the full catalogues. However, in order to reduce the dimensionality of the data vector efficiently, while still retaining most information about the cosmology, similar to previous KiDS-1000 analyses (Loureiro et al., 2022; Tröster et al., 2022), we choose observed angular power spectra, a.k.a. pseudo-Cls (Peebles, 1973), as our data vector of choice. This allows us to compare our SBI analysis to the results from the standard analysis of KiDS-1000 data using pseudo-Cls in Loureiro et al. (2022). The KiDS-SBI pipeline is, in principle, set up to make cosmological inferences from cosmic shear based on any statistic which could be derived from a galaxy catalogue, but the underlying statistical random fields are currently only capable of achieving percent-level accuracy for two-point statistics. Any higher-order statistics measured from the current forward simulations are not necessarily accurate enough to conduct inference. The statistics and their post-processing are defined in Sect. 4.2.
@@
The forward simulations are designed to model weak gravitational lensing observations on the level of galaxy catalogues. The galaxy populations are sampled such that they trace the underlying log-normal random matter fields, while the galaxy shapes are lensed in accordance with the lensing potential of the matter fields along the line-of-sight. As the simulations do not model images, any image-level systematic effects, such as shear biases or variable depth, are included as spatially varying probability density functions.
3.1 Cosmology dependence: three-dimensional matter power spectrum
@@
The basis of the forward simulations is the 3-dimensional matter power spectrum, $P_{\delta,\mathrm{nl}}(k,z;\mathbf{\Theta})$. In the linear regime, the equal-time three-dimensional matter power spectrum’s cosmology dependence comes from the following relation
@@
We compute the non-linear matter power spectrum via a non-perturbative model, HMCode-2016 (Mead et al., 2015, 2016). Although updated iterations of this model exist (Mead et al., 2021), we limit ourselves to HMCode-2016, so our model remains comparable to the KiDS-1000 cosmic shear analysis. HMCode uses a halo model approach to incorporate the effects of baryonic feedback on the matter distribution. The main driving factor in this is the baryonic matter expelled by Active Galactic Nuclei (AGN). This suppresses the power at small scales, i.e. large $k$, as a function of the amplitude of the halo mass-concentration relation, $A_{\mathrm{bary}}$, and the halo bloating parameter, $\eta_{0}$. We also fix the relation between these two parameters to $\eta_{0}=0.98-0.12A_{\mathrm{bary}}$ in accordance with Joudaki et al. (2018) and Joachimi et al. (2021), and treat $A_{\mathrm{bary}}$ as the only free parameter related to baryonic feedback.
We note that this treatment may lead to systematic biases at small scales, as the galaxies will be sampled from a matter power spectrum which is already modified by baryonic feedback. It would be more physically accurate to sample the galaxies from a matter power spectrum without baryonic non-linearities and subsequently add a perturbation to the matter density contrast fields due to baryonic feedback (similar to the approach used for the intrinsic alignments described in Sect. 3.5). Nevertheless, we do not expect this to cause a large discrepancy for the scales which we are probing in the cosmic shear signal ($\ell<1500$). Field-level implementations of baryonic feedback already exist (Schneider & Teyssier, 2015; Schneider et al., 2019, 2022; Porqueres et al., 2023), and they would constitute interesting avenues for future extensions to KiDS-SBI.

### v2

@@ 2 KiDS-1000 data
The Kilo-Degree Survey (KiDS) is a large public galaxy survey conducted by the European Southern Observatory using the OmegaCAM CCD mosaic camera (Kuijken, 2011), attached to the 2.6 m VLT Survey Telescope (VST). The survey covers approximately 1,350 deg2 between two distinct fields known as KiDS North and South, which are fields located in the northern and southern galactic caps (with DEC $\approx$ 0º and DEC $\approx$ -30º, respectively). This area is the same as the one covered by the VISTA Kilo-degree INfrared Galaxy survey (VIKING; Edge et al. 2013), which means that both surveys together observe every object with a total of nine photometric bands: $ugriZYJHK_{\rm s}$. The analysis presented in this paper makes use of DR4 (Kuijken et al., 2019), also known as KiDS-1000, which covers approximately 1,000 deg2. The effective area covered by KiDS-1000 shape measurements (shown in Fig. 1) extends over 773.3 deg2 (Joachimi et al., 2021). This area is calculated directly from the mosaic mask defined a the native OmegaCAM pixel scale of 0.213 arcsec.
@@
The KiDS images were processed using Astro-WISE (McFarland et al., 2013) and the methodology outlined in Wright et al. (2019) to combine the VST and VIKING photometric measurements. To obtain the shape measurements, Theli (Erben et al., 2005) was used to process the $r$-band images, while the gravitational shear estimates are then obtained from lensfit (Miller et al., 2007, 2013; Fenech Conti et al., 2017). Depending on the shape noise variance and the ellipticity measurement noise variance, lensfit assigns a weight to the shear measurement of each galaxy, $w_{i}$, which scales the shear signal such that the signal-to-noise ratio is optimal. Although KiDS-SBI currently does not model lensfit weights (all galaxies are weighted equally by default), the weights are carried through the measurement pipeline when measuring the cosmic shear summary statistics from the real KiDS-1000 data. If the lensfit weight is mischaracterised, this may lead to shear bias. For this purpose, the shape measurements in the KiDS-1000 gold sample are calibrated to determine any shear biases due to selection biases, noise, weight bias, point spread function (PSF) residuals, or otherwise (Giblin et al., 2021). The resulting estimate of the linear multiplicative shear bias per tomographic bin is then used to unbias the galaxy shear measurements in KiDS-1000. All details on the reduction of the images and the shape calibration are given in Kuijken et al. (2019); Wright et al. (2019); Giblin et al. (2021).
@@
The photometric redshifts were estimated from the KiDS 9-band photometry using Bayesian template-fitting as incorporated within the BPZ code (Benítez, 2000; Wright et al., 2019, 2020). The photometric redshifts are calibrated with spectroscopic redshift measurements made of galaxies in the KiDS sample (Hildebrandt et al., 2021) based on a subsample of $99\%$ completeness. The mapping from photometric to spectroscopic redshift, $P(z|z_{\mathrm{ph}})$, is learnt through a self-organising map (Wright et al., 2020; Hildebrandt et al., 2021). As is the case for previous KiDS-1000 analyses, we use the KiDS-1000 ’gold-sample’, which only considers galaxies with reliable redshift and shape measurements. The maximum posterior photometric redshift is subsequently used to split the source galaxy catalogue into five tomographic bins at the following boundaries: $(0.1,0.3,0.5,0.7,0.9,1.2]$. These give the redshift distributions shown in Fig. 2. As the cosmological constraints are sensitive to shifts in the redshift distributions, the KiDS-1000 analysis also allows the shift in the mean of each tomographic bin to vary freely in the cosmological inference as a nuisance parameter while incorporating for the correlations between them (Asgari et al., 2021; Heymans et al., 2021; van den Busch et al., 2022). This not only avoids biases in the cosmological parameter estimates, but also provides an additional test on the redshift calibration.
We also note that KiDS-SBI treats the KiDS-1000 catalogue as a population of point sources unaffected by source blending, since the KiDS data already accounts for overlapping sources in multiple ways. Firstly, the KiDS photometry integrates Gaussian Aperture and PSF photometry (GAaP; Kuijken et al. 2015, 2019; de Jong et al. 2017), where a Gaussian aperture is applied to the images of all sources that is optimised for S/N in the fluxes while also avoiding source blending. An aperture size of 2” is found to be optimal for these purposes and applied to all sources (Kuijken et al., 2015, 2019; de Jong et al., 2017). Secondly, at the source detection step with SExtractor (Bertin & Arnouts, 1996), explicit flagging of overlapping sources occurs. Approximately 5% of sources (Kuijken et al., 2019; Giblin et al., 2021) are excluded from the sample due to this selection. Additionally, lensfit weights also account for any biases due to irregular objects which are not affected by the selection. We note that deblending of sources is not attempted at any point. Lastly, at the calibration stage, the KiDS-1000 shape measurements are validated against image simulations from Kannawadi et al. (2019). The multiplicative shear bias values obtained in this analysis (Giblin et al., 2021) incorporate any residual biases due to correlated noise in the shape measurements, including source blending.
@@ 3 Forward simulations
The forward simulations for cosmic shear analyses described in this paper are fully implemented and available in the KiDS-SBI [^7] Kilo-Degree Survey – Simulation-Based Inference; https://github.com/mwiet/kids_sbi module. This code is built within the Cosmological Survey Inference System (CosmoSIS [^8] https://github.com/joezuntz/cosmosis; Zuntz et al. 2015) and based upon the KiDS Cosmology Analysis Pipeline (KCAP [^9] https://github.com/KiDS-WL/kcap; Joachimi et al. 2021; Asgari et al. 2021; Heymans et al. 2021; Tröster et al. 2021). In addition, the KiDS-SBI pipeline is generalisable to other cosmic shear analyses.
Cosmological parameters, $\mathbf{\Theta}$ 3D matter power spectrum, $P_{\delta,\mathrm{nl}}(\mathbf{k},z)$: Sect. 3.1 Baryonic feedback Split space into concentric shells: $i,j...$: Sect. 3.2 Resolution along the line of sight Non-Limber 2D projection: Sect. 3.3 2D power spectra, $C_{\delta\delta}^{(ij)}(\ell)$ Generate log-normal random fields: Sect. 3.4 Matter, $\delta^{(i)}(\bm{\theta})$, & convergence fields, $\kappa^{(i)}(\bm{\theta})$ Intrinsic alignments, $\kappa_{\mathrm{IA}}^{(i)}(\bm{\theta})$: Sect. 3.5 Sample galaxies: Sect. 3.6 &\&& 3.7 Angular galaxy positions, $\bm{\theta}$, redshifts, $z_{\mathrm{true}}$, observed tomographic bin, $p$, and observed ellipticity, $\epsilon$ Survey characteristics Mask Redshift distributions Shear biases Point spread function Galaxy density Intrinsic galaxy shapes Depth variability: Sect. 3.8 Measure and correct shear: Sect. 4.1 Observed shear, $\epsilon^{\mathrm{corr}}$ Measure pseudo-Cls, subtract noise and binning: Sect. 4.2 Pseudo-Cl, $\tilde{C}_{\epsilon\epsilon}^{(pq)}(\ell)$
@@
The overall outline of the forward simulations is shown in Fig. 3. Firstly, the simulation is fed a set of cosmological and astrophysical parameters, $\mathbf{\Theta}$, from a sampler of our choice. All of these parameters are passed into CAMB (Lewis et al., 2000; Lewis & Challinor, 2002; Howlett et al., 2012) which calculates the 3D matter power spectrum, $P_{\delta,\mathrm{nl}}$, at a given cosmology (see also Sect. 3.1 for more details). We incorporate baryonic feedback using HMCode (Mead et al., 2016). Next, the simulation splits the light cone along the line of sight into concentric spherical shells centred at the observer, as required by our large-scale structure simulator of choice, GLASS (Tessore et al., 2023). Such an approach is computationally efficient as all observables are simulated with a coarse resolution along the line of sight. Within each shell, we sample log-normal random matter fields to efficiently model the two-point statistics with high accuracy, while also incorporating some higher-order fluctuations. As an added benefit, this simplifies the implementation of effects such as redshift evolution as well as survey characteristics, as they usually also follow a spherical geometry. On the other hand, the concentric shells also allow for the Universe to be split into shells, which are of a scale large enough such that a log-normal field becomes an accurate description of the matter over-densities within it (see Sect. 3.2 for details).
@@
Once we know the geometry of our matter shells, we can project the 3D matter power spectrum using the redshift distributions of these shells in order to obtain their 2D angular power spectra, $C_{\delta\delta}$. The latter describes the correlations between the matter over-densities in each shell. Since we also require accurate cross-correlations between different shells to compute the log-normal fields, the Limber approximation is not applicable here, as the cross-correlations are zero in that approximation with the shells being non-overlapping along the line of sight. To avoid this, we project the 3D matter power spectrum with a non-Limber integral. This is usually an expensive computation, so we show here a novel implementation of the Levin method of integration (Levin, 1996), which allows for the efficient integration of the spherical Bessel functions found in non-Limber integrals (more details are given in Sect. 3.3).
@@
We can then construct log-normal matter fields which are consistent with these correlations and their associated cosmology. To do so, we used the Generator for Large Scale Structure (GLASS [^10] Generator for Large Scale Structure; https://github.com/glass-dev/glass, Tessore et al. 2023), which efficiently generates correlated log-normal random matter fields, $\delta$, for each shell. Then, it accurately computes the respective convergence field, $\kappa$, (see Sect. 3.4) while also incorporating an effective convergence field due to intrinsic galaxy alignments, $\kappa_{\mathrm{IA}}$ (see Sect. 3.5).
@@
Upon the construction of the matter, $\delta$, and convergence fields, $\kappa$, we used them to Poisson-sample galaxies that reside within the matter fields as well as their properties, namely, intrinsic ellipticities and shear. During this step, with the aid of the SALMO module (Joachimi et al., 2021), we also consider survey characteristics, such as the survey footprint, along with its spatial variability in depth and in its redshift distributions, as well as in the shape and shot noise, as described in Sect. 3.6. Thus, we obtained a full galaxy catalogue with shape information consistent with the input cosmology as well as with the previously mentioned survey properties.
@@
The last steps of the forward simulation pipeline involve the post-processing and compression of the simulated data catalogue. In principle, any data vector could be obtained here, such as two-point shear statistics (Asgari et al., 2021; van den Busch et al., 2022), the shear fields (Porqueres et al., 2022; Loureiro et al., 2023), or even the full catalogues. However, in order to reduce the dimensionality of the data vector efficiently, while still retaining most information about the cosmology, similar to previous KiDS-1000 analyses (Loureiro et al., 2022; Tröster et al., 2022), we chose the observed angular power spectra, also known as ’pseudo-Cls’ (Peebles, 1973), as our data vector of choice. This allowed us to compare our SBI analysis to the results from the standard analysis of KiDS-1000 data using pseudo-Cls in Loureiro et al. (2022). The KiDS-SBI pipeline is, in principle, set up to make cosmological inferences from cosmic shear based on any statistic which could be derived from a galaxy catalogue, but the underlying statistical random fields are currently only capable of achieving percent-level accuracy for two-point statistics. Any higher-order statistics measured from the current forward simulations are not necessarily accurate enough to conduct inference. The statistics and their post-processing are defined in Sect. 4.2.
@@
The forward simulations are designed to model weak gravitational lensing observations on the level of galaxy catalogues. The galaxy populations are sampled such that they trace the underlying log-normal random matter fields, while the galaxy shapes are lensed in accordance with the lensing potential of the matter fields along the line of sight. As the simulations do not model images, any image-level systematic effects, such as shear biases or variable depth, are included as spatially varying probability density functions.
3.1 Cosmology dependence: 3D matter power spectrum
@@
The basis of the forward simulations is the 3D matter power spectrum, $P_{\delta,\mathrm{nl}}(k,z;\mathbf{\Theta})$. In the linear regime, the equal-time 3D matter power spectrum’s cosmology dependence comes from the following relation:
@@
We can compute the non-linear matter power spectrum via a non-perturbative model, HMCode-2016 (Mead et al., 2015, 2016). Although updated iterations of this model exist (Mead et al., 2021), we limited our study to HMCode-2016, so our model remains comparable to the KiDS-1000 cosmic shear analysis. HMCode uses a halo model approach to incorporate the effects of baryonic feedback on the matter distribution. The main driving factor in this is the baryonic matter expelled by active galactic nuclei (AGNs). This suppresses the power at small scales, namely, large values of $k$, as a function of the amplitude of the halo mass-concentration relation, $A_{\mathrm{bary}}$, and the halo bloating parameter, $\eta_{0}$. We also fixed the relation between these two parameters to $\eta_{0}=0.98-0.12A_{\mathrm{bary}}$ in accordance with Joudaki et al. (2018) and Joachimi et al. (2021), and treated $A_{\mathrm{bary}}$ as the only free parameter related to baryonic feedback.
We note that this treatment may lead to systematic biases at small scales, as the galaxies will be sampled from a matter power spectrum which is already modified by baryonic feedback. It would be more physically accurate to sample the galaxies from a matter power spectrum without baryonic non-linearities and subsequently add a perturbation to the matter density contrast fields due to baryonic feedback (similar to the approach used for the intrinsic alignments described in Sect. 3.5). Nevertheless, we do not expect this to cause a large discrepancy for the scales which we are probing in the cosmic shear signal ($\ell<1500$). Field-level implementations of baryonic feedback already exist (Schneider & Teyssier, 2015; Schneider et al., 2019, 2022; Porqueres et al., 2023) and they would constitute interesting avenues for future extensions to KiDS-SBI.